# From Restless to Contextual: A Thresholding Bandit Approach to Improve Finite-horizon Performance

### v2

@@ Appendix A Whittle’s Index
s.t. $\sum_{m=1}^{M}a_{t}^{m}\leq B,\forall t\in[T].$ (A.9)
We will denote $J^{*}$ to be the optimal value of Problem A.9. Instead of directly solving Problem A.9, Whittle deals with the Lagrangian relaxation of Problem A.9, which is stated in Problem A.10
$\max_{\bm{\pi}_{d}}~{}\liminf_{T\to\infty}\frac{1}{TM}\mathbb{E}\left[\sum_{m=1}^{M}\sum_{t=0}^{T}R^{m}\left(s_{t}^{m}\right)+\lambda\left(1-a_{t+1}^{m}\right)\mid\bm{s}_{0}\right].$ (A.10)
The key observation to derive Whittle’s index policy is that Problem A.10 can be solved by addressing the following dynamic programming separately for each agent
@@
$\beta^{*}_{m}:=\max~{}\liminf_{T\to\infty}\frac{1}{T}\mathbb{E}\left[\sum_{t=0}^{T}R^{m}\left(s_{t}^{m}\right)\mid s_{0}^{m}\right].$
The Whittle’s index $\lambda^{m}_{s}$ is defined as the multiplier $\lambda_{s}^{m}$ when it is equally favorable between choosing action $1$ and action 00, i.e. $Q^{m}_{\lambda^{m}_{s}}(s,1)=Q^{m}_{\lambda^{m}_{s}}(s,0)$. The Whittle’s index policy would then give action to $B$ agents with the highest Whittle’s index at each time step $t$. Denote this as $\bm{\pi}^{\text{Whittle}}$ and the average reward of Whittle’s index policy as $J^{\text{Whittle}}$. The performance of Whittle’s index policy is characterized in Theorem A.1.
Theorem A.1 (Theorem 2, Weber & Weiss 1990).
Suppose all arms are homogeneous, i.e., $P_{a}^{m}=P_{a}^{m^{\prime}},R^{m}=R^{m^{\prime}},\forall m\neq m^{\prime}$, then the Whittle’s index policy is asymptotically optimal,
@@
s.t. $\sum_{m=1}^{M}a_{t}^{m}\leq B,\forall t\in[T].$ (A.11)
Similarly, instead of directly solving Problem A.11, Whittle deals with the Lagrangian relaxation of Problem A.11, which is stated in Problem A.12
$\max_{\bm{\pi}_{d}}~{}\sum_{t=0}^{\infty}\frac{1}{M}\mathbb{E}\left[\sum_{m=1}^{M}\beta^{t}\left(R^{m}\left(s_{t}^{m}\right)+\lambda\left(1-a_{t+1}^{m}\right)\right)\mid\bm{s}_{0}\right].$ (A.12)
Again, the same key observation is that Problem A.12 can be solved by addressing the following dynamic programming separately for each agent
@@
$+\beta\sum_{s^{\prime}\in\mathcal{S}}P^{m}_{a}(s,s^{\prime})V^{m}(s^{\prime}).$ (A.13)
@@
A.1 Generalized Regret in Online Restless Bandits
Given $K$ episodes each with a horizon length of $T$ and the set of online Whittle’s index policy at each episode $k\in[K]$, $\bm{\omega}:=\{\bm{\omega}_{k}\}_{k\in[K]}$, the regret $\mathcal{R}_{f}(\bm{\omega})$:
$\mathcal{R}_{f}(\bm{\omega}):$ $=\frac{1}{K}\sum_{k=1}^{K}\Bigg{(}\mathbb{E}_{\bm{\pi}^{\text{Whittle}}}\left[\sum_{m=1}^{M}\sum_{t=0}^{T}R^{m}\left(s_{t}^{m}\right)\mid\bm{s}_{0}\right]-\mathbb{E}_{\bm{\omega}_{k}}\left[\sum_{m=1}^{M}\sum_{t=0}^{T}R^{m}\left(s_{t}^{m}\right)\mid\bm{s}_{0}\right]\Bigg{)}.$ (A.14)
@@ A.1 Generalized Regret in Online Restless Bandits
Existing regret bounds on online Whittle’s index policy are only meaningful when the number of episodes is sufficiently large. To illustrate this, Wang et al. (2023b) shows that their algorithm has a regret bound of the order $\mathcal{O}\left(T\sqrt{TK}\right)$. If we only have $1$ episode, this will be equivalent to saying that the regret bound is $\mathcal{O}(T^{3/2})$. The hardness lies in estimating the MDP: Azar et al. 2017 show that when $B=M=1$, the lower bound of the regret is $\Omega(T\sqrt{K})$. This lower bound indicates that when directly dealing with an MDP, it is impossible to get a regret bound that is sublinear in $T$.
Appendix B Counter Example for Heterogeneous Agents
In this section, we will construct an instance involving heterogeneous agents where the greedy policy is not optimal. Say that each agent is associated with two state: State 0 and State 1, and the agent can take either Action $1$ (intervention) or Action 00 (no intervention). Assume for now that there are 4 agents with different transition probabilities. The first two agents fall under Type 1: their transition probability under Action 1 is shown in Figure B.4 and their transition probability under Action 0 is shown in Figure B.5. The other two agents fall under Type 2: their transition probability under Action 1 is shown in Figure B.6 and their transition probability under Action 0 is shown in Figure B.7.
@@ Appendix B Counter Example for Heterogeneous Agents
Figure B.4: Dynamic of Agent Type 1 with Intervention
@@
Figure B.5: Dynamic of Agent Type 1 without Intervention
@@
Figure B.6: Dynamic of Agent Type 2 with Intervention
@@
Figure B.7: Dynamic of Agent Type 2 without Intervention
Suppose the budget equals one, meaning that we can only give one agent intervention at each time. We observe that under our problem setup Type 2 agents always have a higher incremental reward $\phi^{2}(1,1):=\frac{3}{4}\cdot 1,\phi^{2}(0,1):=\frac{3}{4}\cdot 1,\phi^{2}(s,0)=0,I^{2}(1)=I^{2}(0)=\frac{3}{4}$, opposed to $\phi^{1}(0,1)=\frac{2}{3}\cdot 1,\phi^{1}(1,1)=\frac{1}{2}\cdot 1,\phi^{1}(0,0)=0,\phi^{1}(1,0)=\frac{1}{3},I(1)=\frac{2}{3},I(0)=\frac{1}{6}$. However, this does not imply that the optimal policy is to always take Action 1 for Type 2 agents. If all resources are allocated to Type 2 agents, the expected reward is $3/4$ per time step summed across all agents. Alternatively, if resources are devoted to Type 1 agents, alternating actions between the two Type 1 agents, the expected reward is $2/3+2/3\times 1/3=8/9$ per Type 1 agent over two time periods (where an action is taken in one period but not in the other). To illustrate this point, we ran the following experiment: we set the horizon to be $20$ and repeated the simulation $300$ times. We compare the performance of the following two policies in terms of cumulative rewards: Policy 1) choosing one of the Type 2 agents for intervention at each time, and Policy 2) alternating intervention between the two Type 1 agents. The observed cumulative for Policies 1 and 2 are 15.02 and 16.78, respectively. This shows that the greedy policy is not optimal under all heterogeneous agents cases even under Assumption 3.1.
Appendix C Technical Details In Section 3
First, recall that at each time $t$ of an offline restless bandit problem, the learner observes the set of states of $M$ agents, denoted by $\bm{s}_{t}:=\left(s_{t}^{1},s_{t}^{2},\cdots,s_{t}^{M}\right)$, and receives the corresponding reward $\bm{r}_{t}:=(r_{t}^{1},\cdots,r_{t}^{M})$. Based on the observed states, the learner determines the action $\bm{a}_{t+1}:=\left(a_{t+1}^{1},a_{t+1}^{2},\cdots,a_{t+1}^{M}\right)$, according to a deterministic policy $\pi^{d}_{t+1}(\bm{s}):\mathcal{S}^{M}\rightarrow\{0,1\}^{M}$. The learner is provided with a budget constraint, allowing at most $B$ arms to be pulled at each time step, i.e., $\sum_{m\in[M]}a_{t+1}^{m}\leq B$. After pulling up to $B$ arms, the learner observes the next state $\bm{s}_{t+1}$, generated according to $s_{t+1}^{m}\sim P_{a_{t+1}^{m}}^{m}\left(s_{t}^{m},s_{t+1}^{m}\right)$, and receives a feedback $\bm{r}_{t+1}$.
To ease notation, in Section C we will use $q^{a}_{ss^{\prime}}(m)$ to denote $P_{a}^{m}(s,s^{\prime})$. Before presenting the proofs, we reformulate the online restless bandit problem as a single large MDP, where the state and action are defined as tuples comprising the composite states and actions of all agents. This reformulation is valid because the agents/arms in the restless bandit problem are independent. This MDP can be described as $\left(\mathcal{S}^{M},\mathcal{F},q,\bm{R}\right)$, where $\mathcal{F}$ is defined as
@@ Appendix C Technical Details In Section 3
$\mathbb{E}_{\bm{\pi}}\left[\sum_{t=h}^{T}\bm{R}(\bm{s}_{t})\mid\bm{s}_{h}=\bm{s}\right].$ (C.15)
@@
Again, we use $\bm{\pi}^{*}:=\{\pi_{t}^{*}\}_{t\in[T]}$ to denote the maximizer, and let $V_{h}^{*}$ be the value function under policy $\bm{\pi}^{*}$. We will use this notation in Appendix C.2, C.3, C.4, and D.
Next, we state the proofs.
C.1 Proof of Lemma 3.2
See 3.2

### v4

@@ Appendix B Whittle’s Index
s.t. $\sum_{m=1}^{M}a_{t}^{m}=B,\forall t\in[T].$ (B.5)
We will denote $J^{*}$ to be the optimal value of Problem B.5. Instead of directly solving Problem B.5, Whittle deals with the Lagrangian relaxation of Problem B.5, which is stated in Problem B.6
$\max_{\bm{\pi}_{d}}~\liminf_{T\to\infty}\frac{1}{TM}\mathbb{E}\left[\sum_{m=1}^{M}\sum_{t=0}^{T}R^{m}\left(s_{t}^{m}\right)+\lambda\left(1-a_{t+1}^{m}\right)\mid\bm{s}_{0}\right].$ (B.6)
The key observation to derive Whittle’s index policy is that Problem B.6 can be solved by addressing the following dynamic programming separately for each agent
@@
$\beta^{*}_{m}:=\max~\liminf_{T\to\infty}\frac{1}{T}\mathbb{E}\left[\sum_{t=0}^{T}R^{m}\left(s_{t}^{m}\right)-\lambda\mathbbm{1}\{a=1\}\mid s_{0}^{m}\right].$
The Whittle’s index $\lambda^{m}_{s}$ is defined as the multiplier $\lambda_{s}^{m}$ when it is equally favorable between choosing action $1$ and action 0, i.e. $Q^{m}_{\lambda^{m}_{s}}(s,1)=Q^{m}_{\lambda^{m}_{s}}(s,0)$. The Whittle’s index policy would then give action to $B$ agents with the highest Whittle’s index at each time step $t$. Denote this as $\bm{\pi}^{\text{Whittle}}$ and the average reward of Whittle’s index policy as $J^{\text{Whittle}}$. The performance of Whittle’s index policy is characterized in Theorem B.1.
Theorem B.1 (Theorem 2, Weber & Weiss 1990).
Suppose all arms are homogeneous and assumptions therein, i.e., $P_{a}^{m}=P_{a}^{m^{\prime}},R^{m}=R^{m^{\prime}},\forall m\neq m^{\prime}$, then the Whittle’s index policy is asymptotically optimal,
@@
s.t. $\sum_{m=1}^{M}a_{t}^{m}=B,\forall t\in[T].$ (B.7)
Similarly, instead of directly solving Problem B.7, Whittle deals with the Lagrangian relaxation of Problem B.7, which is stated in Problem B.8
$\max_{\bm{\pi}_{d}}~\sum_{t=0}^{\infty}\frac{1}{M}\mathbb{E}\left[\sum_{m=1}^{M}\beta^{t}\left(R^{m}\left(s_{t}^{m}\right)+\lambda\left(1-a_{t+1}^{m}\right)\right)\mid\bm{s}_{0}\right].$ (B.8)
Again, the same key observation is that Problem B.8 can be solved by addressing the following dynamic programming separately for each agent
@@
$+\beta\sum_{s^{\prime}\in\mathcal{S}}P^{m}_{a}(s,s^{\prime})V^{m}(s^{\prime}).$ (B.9)
@@
B.1 Generalized Regret in Online Restless Bandits
Given $K$ episodes each with a horizon length of $T$ and the set of online Whittle’s index policy at each episode $k\in[K]$, $\bm{\omega}:=\{\bm{\omega}_{k}\}_{k\in[K]}$, the regret $\mathcal{R}_{\text{Restless}}^{\bm{\pi}^{\text{Whittle}}}(\bm{\omega})$:
$\mathcal{R}_{f}(\bm{\omega}):$ $=\frac{1}{K}\sum_{k=1}^{K}\Bigg(\mathbb{E}_{\bm{\pi}^{\text{Whittle}}}\left[\sum_{m=1}^{M}\sum_{t=0}^{T}R^{m}\left(s_{t}^{m}\right)\mid\bm{s}_{0}\right]-\mathbb{E}_{\bm{\omega}_{k}}\left[\sum_{m=1}^{M}\sum_{t=0}^{T}R^{m}\left(s_{t}^{m}\right)\mid\bm{s}_{0}\right]\Bigg).$ (B.10)
@@ B.1 Generalized Regret in Online Restless Bandits
Existing regret bounds on online Whittle’s index policy are only meaningful when the number of episodes is sufficiently large. To illustrate this, Wang et al. (2023b) shows that their algorithm has a regret bound of the order $\mathcal{O}\left(T\sqrt{K}\right)$. If we only have $1$ episode, this will be equivalent to saying that the regret bound is $\mathcal{O}(T)$. The hardness lies in estimating the MDP: Auer et al. 2008 show that the lower bound of the regret is $\Omega(T\sqrt{K})$. This lower bound indicates that when directly dealing with an MDP, it is impossible to get a regret bound that is sublinear in $T$.
Appendix C Counter Example for Heterogeneous Agents
In this section, we will construct an instance involving heterogeneous agents where the greedy policy is not optimal. Say that each agent is associated with two state: State 0 and State 1, and the agent can take either Action $1$ (intervention) or Action 0 (no intervention). Assume for now that there are 4 agents with different transition probabilities. The first two agents fall under Type 1: their transition probability under Action 1 is shown in Figure C.2 and their transition probability under Action 0 is shown in Figure C.3. The other two agents fall under Type 2: their transition probability under Action 1 is shown in Figure C.4 and their transition probability under Action 0 is shown in Figure C.5.
@@ Appendix C Counter Example for Heterogeneous Agents
Figure C.2: Dynamic of Agent Type 1 with Intervention
@@
Figure C.3: Dynamic of Agent Type 1 without Intervention
@@
Figure C.4: Dynamic of Agent Type 2 with Intervention
011p=1p=1p=0p=0p=0p=0p=1p=1r=0r=0r=1r=1
Figure C.5: Dynamic of Agent Type 2 without Intervention
Suppose the budget equals one, meaning that we can only give one agent intervention at each time. We observe that under our problem setup Type 2 agents always have a higher incremental reward $\phi^{2}(1,1):=\frac{3}{4}\cdot 1,\phi^{2}(0,1):=\frac{3}{4}\cdot 1,\phi^{2}(s,0)=0,I^{2}(1)=I^{2}(0)=\frac{3}{4}$, opposed to $\phi^{1}(0,1)=\frac{2}{3}\cdot 1,\phi^{1}(1,1)=\frac{1}{2}\cdot 1,\phi^{1}(0,0)=0,\phi^{1}(1,0)=\frac{1}{3},I(1)=\frac{2}{3},I(0)=\frac{1}{6}$. However, this does not imply that the optimal policy is to always take Action 1 for Type 2 agents. If all resources are allocated to Type 2 agents, the expected reward is $3/4$ per time step summed across all agents. Alternatively, if resources are devoted to Type 1 agents, alternating actions between the two Type 1 agents, the expected reward is $2/3+2/3\times 1/3=8/9$ per Type 1 agent over two time periods (where an action is taken in one period but not in the other). To illustrate this point, we ran the following experiment: we set the horizon to be $20$ and repeated the simulation $300$ times. We compare the performance of the following two policies in terms of cumulative rewards: Policy 1) choosing one of the Type 2 agents for intervention at each time, and Policy 2) alternating intervention between the two Type 1 agents. The observed cumulative for Policies 1 and 2 are 15.02 and 16.78, respectively. This shows that the greedy policy is not optimal under all heterogeneous agents cases.
Appendix D Counter Example for Three-State, Homogeneous Agents
In this section, we will construct an instance involving homogeneous agents with three states where the greedy policy is not optimal. Say that each agent is associated with three states: State 0, State 1 and State 2, and the agent can take either Action 1 (intervention) or Action 0 (no intervention). Assume for now that there are two agents with the same transition probabilities. Their transition probability under Action 1 is shown in Figure D.6 and their transition probability under Action 0 is shown in Figure D.7.
@@ Appendix D Counter Example for Three-State, Homogeneous Agents
Figure D.6: Dynamic with Intervention
@@
Figure D.7: Dynamic without Intervention
Suppose that the budget equals one, we observe that State 2 and State 1 have the same incremental reward $I(2)=\frac{3}{7}-\frac{5}{28}=\frac{1}{4}=\frac{3}{2}-\frac{5}{4}=I(1)$. However, we will show that when $T=3$, greedy policy is not optimal. Specifically, we will show that when $t=0$ and $\bm{s}_{0}=(1,2)$, the optimal policy would be giving State 2 Action 1 instead of State 1 Action 1. This will show that greedy policy is not optimal as greedy policy could give State 1 Action 1 by definition. Suppose we are giving State 1 Action 1, then we have
$V((1,2))$ $=3+\frac{5}{56}V_{1}^{*}((1,2))+\frac{5}{56}V_{1}^{*}((1,1))+\frac{23}{28}V_{1}^{*}((1,0))$
$=3+\frac{5}{56}\frac{47+3*28}{28}+\frac{5}{56}\frac{19}{4}+\frac{23}{28}\frac{5}{2}.$
Similarly, if we are giving State 2 Action 1, then we have
$V((1,2))$ $=3+\frac{1}{28}V_{1}^{*}((2,2))+\frac{1}{7}V_{1}^{*}((1,2))+\frac{15}{28}V_{1}^{*}((1,0))+\frac{5}{28}V_{1}^{*}((2,0))+\frac{3}{28}V_{1}^{*}((1,1))$
$=3+\frac{1}{28}\left(\frac{17}{28}+4\right)+\frac{1}{7}\left(3+\frac{47}{28}\right)+\frac{15}{28}\left(\frac{5}{2}\right)+\frac{5}{28}\left(\frac{3}{7}+2\right)+\frac{3}{28}\left(\frac{11}{4}+2\right).$
By calculation, we will have the results. This shows that the greedy policy is not optimal under three states setting.
Appendix E Technical Details In Section 3
First, recall that at each time $t$ of an offline restless bandit problem, the learner observes the set of states of $M$ agents, denoted by $\bm{s}_{t}:=\left(s_{t}^{1},s_{t}^{2},\cdots,s_{t}^{M}\right)$, and receives the corresponding reward $\bm{r}_{t}:=(r_{t}^{1},\cdots,r_{t}^{M})$. Based on the observed states, the learner determines the action $\bm{a}_{t+1}:=\left(a_{t+1}^{1},a_{t+1}^{2},\cdots,a_{t+1}^{M}\right)$, according to a deterministic policy $\pi^{d}_{t+1}(\bm{s}):\mathcal{S}^{M}\rightarrow\{0,1\}^{M}$. The learner is provided with a budget constraint, allowing at most $B$ arms to be pulled at each time step, i.e., $\sum_{m\in[M]}a_{t+1}^{m}=B$. After pulling up to $B$ arms, the learner observes the next state $\bm{s}_{t+1}$, generated according to $s_{t+1}^{m}\sim P_{a_{t+1}^{m}}^{m}\left(s_{t}^{m},s_{t+1}^{m}\right)$, and receives a feedback $\bm{r}_{t+1}$.
To ease notation, in Section E we will use $q^{a}_{ss^{\prime}}(m)$ to denote $P_{a}^{m}(s,s^{\prime})$. Before presenting the proofs, we reformulate the online restless bandit problem as a single large MDP, where the state and action are defined as tuples comprising the composite states and actions of all agents. This reformulation is valid because the agents/arms in the restless bandit problem are independent. This MDP can be described as $\left(\mathcal{S}^{M},\mathcal{F},q,\bm{R}\right)$, where $\mathcal{F}$ is defined as
$\mathcal{F}:=\bigg\{$ $\left(a^{1},a^{2},\cdots,a^{M}\right):\left(a^{1},a^{2},\cdots,a^{M}\right)\in\mathcal{A}^{M},\sum_{m=1}^{M}a^{m}\leq B\bigg\}.$
@@ Appendix E Technical Details In Section 3
$\mathbb{E}_{\bm{\pi}}\left[\sum_{t=h}^{T}\bm{R}(\bm{s}_{t})\mid\bm{s}_{h}=\bm{s}\right].$ (E.11)
@@
Again, we use $\bm{\pi}^{*}:=\{\pi_{t}^{*}\}_{t\in[T]}$ to denote the maximizer, and let $V_{h}^{*}$ be the value function under policy $\bm{\pi}^{*}$. We will use this notation in Appendix  E.1, and F.
E.1 Proof of Theorem 3.1
See 3.1